OpenStudy (nuccioreggie):

Find negative square root of 16. −4 4 ±4 −8

OpenStudy (nuccioreggie):

I think its -4

OpenStudy (anonymous):

So find \[\large -\sqrt{16}~?\]

OpenStudy (nuccioreggie):

-4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Is that what you're asking?

OpenStudy (nuccioreggie):

Yes im asking am i correct

OpenStudy (anonymous):

im sorry I wasn't paying attention the correct answer is -4

OpenStudy (anonymous):

Well it doesn't have a real answer. It ends up as \[\large -4*-4=16\] not \[\large -4*-4=-16\] The only way to solve this is to add on an imaginary number i.

OpenStudy (nuccioreggie):

Yea so its A correct?

OpenStudy (anonymous):

Well what is \[\large -4^{2}~?\]

OpenStudy (mrnood):

it is not \[\sqrt{-16}\] it is \[-\sqrt{16}\] (well even this is not quite true sqrt 16 is +4 or -4 the negative of those is obvious BUT -sqrt(16) is still +4 OR -4

OpenStudy (mrnood):

@LegendarySadist there is no need for imaginary solutions here